Give examples where (a) the velocity of a particle is zero but its acceleration is not zero, (b) the velocity is in opposite direction to the acceleration , (c) the velocity is perpendicular to the acceleration.

Grade:12

2 Answers

vikas askiitian expert
13 years ago

a)let particle velocity is given by v=3-t then a=dv/dt=3

at t=3s  ,v=0 but a is 3

b)at t=4s

 v=-ve (backward)

 a=3(forward)

c)in this case we can take example of uniform circular motion in which a particleis moving with a constant velocity in a circle of radius r and the so there exists a

centripital accleration directed towards the center of circle or perpendicular to velocity.

sivaji naidu
13 years ago

(a) suppose a body thrown vertically up and at the maximum point, the velocity becomes 0 but the acceleration is downward with an acceleration g.

(b) in the same case, when the body is moving upward, the velocity is in one direction but the acceleration is towards down (g)